100 Watt 200 volt find current and resistance
1
Expert's answer
2019-11-21T10:03:36-0500

Let current = I ,

Resistance = R


Voltage V = 200 volt


Power ,. W = I V

Or. 100 = I x 200

Or, I = 100 / 200 = 0.5 Amp


Now , V = I R

So,. 200 = 0.5 x R

So,. R = 200 / 0.5 = 400 ohm